OpenStudy (anonymous):

find dy/dx if y=sin(x^2)

OpenStudy (anonymous):

it is 2x cos(x^2)

OpenStudy (anonymous):

\[\huge y=\sin(x^2) \rightarrow \frac{dy}{dx} = \frac{d}{dx} \sin (x^2)\] \[\huge \rightarrow \frac{dy}{dx} = \cos (x^2) \frac{d}{dx} (x^2)\] \[\huge \rightarrow \frac{dy}{dx} = \cos (x^2) 2x\] \[\huge \rightarrow \frac{dy}{dx} =2x \cos (x^2) \]

OpenStudy (anonymous):

@slpisu